OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 5| + 2 = 5 Step 2: |x – 5| = 5 – 2 Step 3: |x – 5| = 3 Which of the following is a correct next step to solve the equation?

OpenStudy (anonymous):

x + 5 = -3 -x – 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (radar):

Personally my next step would go ahead and solve this equation which would be as follows: x - 5 = 3, x=8 Then i would solve for the other value (which I believe your choices are getting at. The absolute value of |x-5| will also have to equal -3, so the equation would be set up to solve for x for that condition.

OpenStudy (anonymous):

Thanks.. :)